1.      Consider in perfectly competitive market the following demand and supply equations for sugar:

Qd =1000-1000p        where Q d is quantity demanded and Qs is quantity supplied. Qs=800+ 1000p

Where P is the price of sugar per pound and Q is thousands of pounds of sugar.

a.  What are the equilibrium price and quantity for sugar? (1point)

b.  Suppose that the government wishes to subsidize sugar production by placing a floor on sugar prices of $0.20 per pound. What would be the relationship between the quantity supplied and quantity demand for sugar?

C. Identify market problem specifically at prices 0.2 per pound and what will be scientific recommendation you suggest to solve the identified market problem?

 


1
Expert's answer
2021-04-30T10:36:37-0400

a:

"Qd=Qs"

"1000-1000p=800+1000P"

"1000-800=1000p+1000p"

"200=2000p"

"P=0.1"

"Q = 800 + 1000(0.1)"

"= 800 + 100"

"Q=900"


b:

"Qs = 800 + 1000(p + 0.2)"

"Qs = 800 + 1000p + 200"

"Qs = 1000 + 1000p"  

The quantity demanded function will remain same: 

"Qd = 1000 - 1000p"


c:

The calculation for quantity demanded and quantity supplied of sugar at $0.2 per pound is:

"Qd=1000-1000p"

"=1000-1000(0.2)"

"=800"

"Qd=800+1000p"

"=800+1000(0.2)"

"=1000"

In the market, the quantity supplied is more than the quantity demanded of sugar at $0.2 per pound. This shows that the seller is not able to sell all the quantity of sugar he manufactures. Due to this, the producer needs to decrease the prices of sugar to increase the demand by making the sugar affordable and cheaper. Now, the increased demand will balance the market by consuming the excess supply that the producer manufacture. Hence, the individual should recommend reducing the prices of sugar to attract more customers in order to increase demand.